OpenStudy (sydthekid913):

find the area of a triangle who's base is x^2+2x+4 and who's height is 2x^2+2x+6 (will fan and medal)

OpenStudy (chosenmatt):

area = 1/2 * base * height

OpenStudy (chosenmatt):

did that help?

OpenStudy (sydthekid913):

I already knew that but thank you anyway \(\color{lime}{\Huge\ddot\smile}\)

OpenStudy (sydthekid913):

@haleyelizabeth2017 @Bookworm14

OpenStudy (triciaal):

so you know the formula to find the area of the triangle substitute the expressions and simplify. use what you know to find the unknown

OpenStudy (triciaal):

find the Area 1/2 * base * height 1/2 * (x^2+2x+4) * (2x^2+2x+6)

OpenStudy (sydthekid913):

I tried but it didn't work :(

OpenStudy (sydthekid913):

I got 2x^3+6x^2+22x+12

OpenStudy (triciaal):

(x +2)(x^2 + x + 3)

OpenStudy (triciaal):

x^3 + x^2 + 3 x + 2 x^2 + 2 x + 6 Area = x^3 + 3 x^2 + 5 x + 6

OpenStudy (triciaal):

check your term in x and did you divide by 2 ?

OpenStudy (sydthekid913):

yes

OpenStudy (sydthekid913):

I tried two times but I still couldn't get it :(

OpenStudy (triciaal):

is my answer correct? do you want to post what you did? the difference between mine and yours is the term in x and the factor

OpenStudy (sydthekid913):

sorry, your answer wasn't one of my choices

OpenStudy (triciaal):

i have an error (x +2)(x^2 + x + 3) should be (x +2)^2

OpenStudy (triciaal):

Area =(x^3 + 3 x^2 + 5 x + 6 )(x + 2) = x^4 + 5 x ^3 + 11 x ^2 + 16 x + 12

OpenStudy (sydthekid913):

sorry, my choices are A. 3x^4+9x^2+10x+12 B. x^4 +3x^3+9x^2+10x+12 C. 3x^4+3x^3+9x^2+10x+12 D. x^4+9x^2+10x+12

OpenStudy (triciaal):

x^4 + 3 x^3 + 9 x^2 + 10 x + 12

OpenStudy (triciaal):

the correct answer is B

OpenStudy (triciaal):

re do the process is correct 1/2 * 2(x^2 + x + 3)(x^2 + 2 x + 4)

OpenStudy (sydthekid913):

thank you :)

OpenStudy (triciaal):

sorry about my previous calculation errors you are welcome